# Attentive AV-FusionNet: Audio-Visual Quality Prediction with Hybrid Attention

### v1

@@ 3 Assessing Modality Importance
$I^{\text{final}}_{m}=\alpha\cdot\text{norm}(I^{\text{abl}}_{m})+\beta\cdot[1-\mathrm{norm}\!\big(I^{\text{norm}}_{m}\big)],\qquad\alpha>\beta,$ (8)
with $\text{norm}(\cdot)$ denoting min–max normalization. Ablation sensitivity is weighted more heavily to reflect direct performance impact, yielding an interpretable measure of modality importance. As shown in Fig.2, the model adapts content-dependently: Fig. 2(a) shows a scene with audio containing dense transients (a known codec challenge [26]), whereas Fig. 2(b) depicts a visually dynamic scene with easy-to-code music. The model appears to assign higher relevance to audio when impairments are salient and to video when visual distractions may mask them. Although subjective data on modality relevance are unavailable, preventing absolute accuracy quantification, observed trends remain consistent when a modality’s bitrate is fixed and the other is varied.
@@ 5.1 Experimental Comparison
Support Vector Regression (SVR). SVR was applied as a regression-based fusion method with Radial Basis Function kernels and multiple feature configurations. Three feature families were tested: (i) $Q_{a}$ and $Q_{v}$ only (SVR-2F), (ii) $Q_{a}$ and $Q_{v}$ plus audio bitrate (SVR-3F), and (iii) extended video features derived from VMAF. Within the latter, two variants were considered: SVR-7F, which includes $Q_{a}$ together with six VMAF features, and SVR-8F, which additionally includes audio bitrate. Video bitrate was intentionally excluded, since the internal dataset provides constant bitrates (CBR), while the external dataset is encoded using a CRF, leading to incompatible feature alignment. On LIVE-SJTU, SVR achieves correlations up to 0.96 and $\operatorname{RMSE}$ as low as 0.17. Performance is lower on the internal dataset, where SVR-8F reaches $R_{p}=0.90$ and $\operatorname{RMSE}=0.43$. Cross-dataset evaluation indicates that models trained on the internal dataset generalize well to LIVE-SJTU, whereas the reverse is less effective, suggesting that LIVE-SJTU has simpler degradation patterns.

### v2

@@ 3 Assessing Modality Importance
$I^{\text{final}}_{m}=\alpha\cdot\mathrm{norm}(I^{\text{abl}}_{m})+\beta\cdot[1-\mathrm{norm}\big(I^{\text{norm}}_{m}\big)],$ (8)
where $\mathrm{norm}(\cdot)$ denotes min–max normalization. We set $\alpha=0.7$ and $\beta=0.3$ to prioritize ablation sensitivity, yielding an interpretable measure of modality importance. As shown in Fig.2, the model adapts content-dependently: Fig. 2(a) shows a scene with audio containing dense transients (a known codec challenge [2]), whereas Fig. 2(b) depicts a visually dynamic scene with easy-to-code music. The model appears to assign higher relevance to audio when impairments are salient and to video when visual distractions may mask them. Although subjective data on modality relevance are unavailable, preventing absolute accuracy quantification, observed trends remain consistent when a modality’s bitrate is fixed and the other is varied.
@@ 5.1 Experimental Comparison
Support Vector Regression (SVR). SVR was applied as a regression-based fusion method with Radial Basis Function kernels and multiple feature configurations. Three feature families were tested: (i) $Q_{a}$ and $Q_{v}$ only (SVR-2F), (ii) $Q_{a}$ and $Q_{v}$ plus audio bitrate (SVR-3F), and (iii) extended video features derived from VMAF. Within the latter, two variants were considered: SVR-7F, which includes $Q_{a}$ together with six VMAF features, and SVR-8F, which additionally includes audio bitrate. Video bitrate was intentionally excluded, since the internal dataset provides constant bitrates (CBR), while the external dataset is encoded using a CRF, leading to incompatible feature alignment. On LIVE-SJTU, SVR achieves correlations up to 0.95 and $\operatorname{RMSE}$ as low as 0.47. Performance is lower on the internal dataset, where SVR-8F reaches $R_{p}=0.90$ and $\operatorname{RMSE}=0.43$. Cross-dataset evaluation indicates that models trained on the internal dataset generalize well to LIVE-SJTU, whereas the reverse is less effective, suggesting that LIVE-SJTU has simpler degradation patterns.